A person stands on a scale in an elevator. What is the elevator's acceleration when the scale reads 3/4 of her weight?
Answer #1

N - Mg = - Ma Hence N= Mg ( 1- a/ g )

here in this case N= 3/4 Mg

(3/4) Mg = Mg ( 1- a/g )

a= g/ 4 = 2.45 m/s2 downwards